OpenStudy (anonymous):

Find (f^-1)'(a) for Sqrt(x^3+x^2+x+1), a=2

OpenStudy (anonymous):

\[find (f^-1)'(a) for \sqrt{x ^{3}+ x^2+x+1} \]

OpenStudy (anonymous):

where a=2

OpenStudy (anonymous):

my guess is \(f^{-1}(2)=1\)

OpenStudy (anonymous):

what inverse equation did you get for that?

OpenStudy (anonymous):

i set \[\sqrt{x^3+x^2+x+1}=2\] and made an educated guess as to what \(x\) should be

OpenStudy (anonymous):

since \(f(1)=\sqrt{4}=2\) that tells you \(f^{-1}(2)=1\)

OpenStudy (anonymous):

Makes sense

OpenStudy (anonymous):

oh i wasn't looking carefully

OpenStudy (anonymous):

But i feel theres a more effective way to prove this

OpenStudy (anonymous):

you don't want \(f^{-1}(2)\) you want the derivative at \(2\)

OpenStudy (anonymous):

the derivative of the inverse, correct?

OpenStudy (anonymous):

no matter, \[f^{-1}(x)=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

in this case \[f^{-1}(2)=\frac{1}{f;(f^{-1}(2))}=\frac{1}{f'(1)}\]

OpenStudy (anonymous):

so your job is 1) find the derivative of \(f(x)=\sqrt{x^3+x^2+x+1}\) 2) find \(f'(1)\) 3) take the reciprocal

OpenStudy (anonymous):

I suppose my confusion rests in how the inverse function of f=1, or did that come from setting it equal to A?

OpenStudy (anonymous):

you can almost do it in your head \[f(x)=\sqrt{x^3+x^2+x+1}\] \[f'(x)=\frac{3x^2+2x+1}{2\sqrt{x^3+x^2+x+1}}\] \[f'(1)=\frac{6}{4}=\frac{3}{2}\] and the reciprocal of \(\frac{3}{2}\) is \(\frac{2}{3}\)

OpenStudy (anonymous):

how did if find \(f^{-1}(2)\)? i think it is called "by inspection" don't try using algebra just think

OpenStudy (anonymous):

ah, all right- i had the order of operations mixed up it seems